((5.15)(0.372))/0.085 =
1
Expert's answer
2020-12-28T04:40:16-0500

((5.15)(0.372))/0.085 =

= "\frac{(5.15)(0.372)}{0.085}"


= "\frac{1.9158}{0.085}"


= 22.5388235294

= 22.54 approx.